OpenStudy (anonymous):

(f^-1) of f(x) = ( x^(3)+2) at (1 , 3).

sam (.sam.):

Inverse of function?

OpenStudy (anonymous):

Yes, please

OpenStudy (anonymous):

the 'at (1, 3)' is throwing me off

OpenStudy (anonymous):

i think its irrelevant? so just find inverse

sam (.sam.):

I think its just a saying that you should let y=f(x), you see, \[f(x)=x^3+2, \] when y=f(x), \[y=x^3+2\] 3=1+2 3=3

OpenStudy (anonymous):

why would it say \(f^{-1}\)? i think its asking for inverse

sam (.sam.):

Inverse goes with \[y=f(x) \\ \\ y=x^3+2 \\ \\ ^3\sqrt{y-2}=x \\ \\ f^{-1}(x)=^3\sqrt{x-2}\]

OpenStudy (anonymous):

Yup, inverse

OpenStudy (anonymous):

yeah correct btw proper way to write cube root is \sqrt[3]{x} \[\sqrt[3]{x}\]

OpenStudy (anonymous):

I'm liking Sam's Answer.

sam (.sam.):

yes with the \(\pm\)

OpenStudy (anonymous):

no you dont need +- because its cube its odd number

sam (.sam.):

\text{} doesn't work btw

OpenStudy (anonymous):

im not using \text{}?

OpenStudy (anonymous):

if you don't believe me that you don't need +- here's a solution generated by computer http://www.wolframalpha.com/input/?i=inverse+of+f%28x%29+%3D+x%5E3%2B2

OpenStudy (anonymous):

I do guys, thanks a lot. I already love this site